# Machine learning for advancing low-temperature plasma modeling and simulation

### v1

@@ 2.3 Plasma-surface interaction
Krüger et al.[123] have conceptualized a PSI ML surrogate model for the Ar ion bombardment of a TiAl composite target as indicated in Figure 3. The ML model, specifically an MLP, was meant to map the incident ions’ energy distributions to the energy angle distributions (EADs) of sputtered particles. It was trained with noisy data from TRIDYN simulations (poor statistics), but demonstrated to be capable of predicting smooth EADs, which were almost congruent to reference TRIDYN simulations with high statistical quality. Gergs et al.[124] have continued those efforts by considering TiAl composite targets with varying compositions, introducing the stoichiometry as a basic system state to the model. A similar data aggregation scheme based on TRIDYN simulations was executed. Though the ML model was replaced by a combination of a convolutional -variational autoencoder (-VAE)[125, 126, 127] and CNN, defining a reduced dimensional (latent) space for the sputtered particles’ EADs and mapping of the incident ion energy distributions as well as TiAl target composition to this particular latent space, respectively. Predictions and reference TRIDYN simulations agreed with each other, even though the assembled PSI ML surrogate model had got merely 0.39 % of the MLP’s[123] degrees of freedom and, hence, a significantly reduced likelihood for overfitting. Kim et al.[128] have pursued a similar but more advanced approach for the sputtering of fluorinated Si surfaces due to Ar ion bombardment. MD instead of TRIDYN simulations were conducted to obtain information on the emitted (i.e., sputtered, reflected) particles’ EADs as a function of the surface F coverage as well as Ar ions’ incident angle (but not kinetic energy). The aggregated EADs were used for a deep learning-based ROM (DL-ROM), comparing the performance of different ML architectures (i.e., autoencoder[27] (AE), denoising AE[129], VAE, conditional VAE (CVAE)[130]) with each other. The CVAE was found to best its competitors. The PSI ML surrogate model was completed by introducing an additional MLP, which mapped the input information (i.e., surface F coverage, Ar ions’ incident angle) to the latent space established by the DL-ROM, enabling resource efficient predictions with atomic, MD fidelity. Gergs et al.[131] have performed hybrid MD/time-stamped force-bias Monte Carlo (tfMC)[132, 133, 134] simulations to setup a PSI ML surrogate model but treated sputtering of Al targets due to Ar ion bombardment and growth of Al thin films at the substrate as generalized wall interactions. The incident particle flux composition as well as Ar ion energy was varied to setup the data set.[135] Two CVAEs were stacked to setup a physics-separating artificial neural network (PSNN), separating the ion bombardment induced damage from the damage to surface state (e.g., composition, mass density) translation. This network architecture has been argued to allow for training with experimental and simulation data in spite of an inherent discrepancy regarding the information accessibility (damage cannot be measured by similar means). The predictions mitigated the noise but maintained key physical features (e.g., threshold for Ar clustering, reflection probability) for small particle doses. In a followup work, [136] have shared a methodology that enabled predictions for particle emission as well as growth during the reactive sputter deposition of AlN thin films in Ar/N plasmas on experimental timescales, that is minutes to hours. First, the MD interaction potential was revised with an evolution strategy implemented in the genetic algorithm-based reactive force field optimizer method (GARFfield)[137] to guarantee accurate PSI simulations.[138] Second, a high-throughput, randomized data generation scheme was pursued to efficiently populate the relevant parameter space with hybrid MD/tfMC simulations.[136] Third, two CVAE were trained to predict either the PSIs or the diffusion processes. They were combined to form a PSNN[131] as PSI ML surrogate model, which can readily be used to complement either plasma simulations (e.g., PIC) or experimental diagnostics. The latter scenario was considered to successfully validate the model. Process times of 45 minutes were predicted within merely 34 GPU hours. In contrast, hybrid MD/tfMC simulations were argued to take more than approximately 8 million CPU years to finish on a comparable case study.
@@ 2.4 Plasma process control and optimization
In a series of contributions, Woelfel et al. have developed MPC schemes for radio-frequency reactive magnetron sputtering fundamentally based on the Berg model.[139] Following the identification of a ROM taking into account the nonlinear process dynamics due to surface poisoning,[140] they have devised a controller for process operation at transition between poisoned and clean surface modes using the previous reduced model, combined with measurements from MRP experiments.[141, 142] An improved reduced model and novel control scheme (also based on the Berg model) has been introduced later, taking into account the plasma state and its dynamical behavior.[143]
@@
The optimization of a plasma etch process has been investigated by Xiao et al. using a MPC scheme based on a physical multi-scale model pairing a 2D fluid description of the discharge and a kinetic Monte Carlo simulation of the surface kinetics. Initially, based on the simulated fluxes at the substrate surface, a RNN surrogate model was set up. Another one was implemented for the surface description, taking into account the stochasticity of the MC model.[59] Subsequently, a similar physical modeling approach has been proposed for a hierarchical surrogate model strategy.[60] Using similar 2D plasma fluid simulation results as for the previous approach, a ROM was devised via proper orthogonal decomposition and Galerkin’s method.[147] This reduced model was argued to capture the intrinsic system dynamics on a low-dimensional representation. It thereafter was used for training another RNN surrogate model for rapid prediction of the plasma discharge dynamics during argon-chlorine etching of silicon. In both works by Xiao et al. MPC schemes have further been developed to control the etching depth and the feature bottom roughness of the process.
@@
With the goal to estimate the internal state of a LTP system in real time despite limited access to observable quantities, a physically-constrained extended Kalman filter (PC-EKF) has been devised by Greve et al.[148] Firstly, the general principle including physical constraints (e.g., prohibiting negative quantities) was assessed at the example of the Lorenz system. The PC-EKF has been shown to closely resemble its chaotic system dynamics. Thereafter, the dynamics of a Hall effect thruster described by a global ionization model were considered. Based on particle balance equations for electrons and ions, the electron temperature was dynamically estimated based on the measured discharge currents. Successful recovery of the electron temperature was demonstrated by the investigation of four different discharge modes subject to varying types of ionization oscillations. The proposed scheme has been argued to benefit from improved robustness with the ability to estimate low-frequency dynamics in the presence of possibly not well resolved high-frequency dynamics (in the model or experiments). In a followup study also by Greve et al.,[149] the PC-EKF approach was generalized to more complex Ar and Ar/O chemical reaction sets and global models of ICP sources. Particle balance equations as well as a power balance equation for electrons were taken into account. In an extensive case study, the robustness of the scheme has been demonstrated. Among other cases, a real time estimate of one or multiple unknown state variables based on corresponding limited measurements (e.g., electron absorbed input power based on argon ion density) has been shown, including pulsed operation.
@@ 3 Advances
The need for LTP and plasma chemistry databases have been previously highlighted (see section 2 and referenced literature).[9, 10, 63] Many data-driven advances in modeling and simulation of LTP science and technology may be enabled only by a strategic data generation initiative. While there is enormous potential in such data generation and exploitation, we reason that this is a matter of scientific practice and awareness but not a technical obstacle, to a substantial extent. In the following, in contrast, we focus on technical aspects that are hypothesized to have the potential to advance the field, proceeding from ML methods to LTP disciplines. We initially discuss ML concepts and methodology of general relevance to LTP modeling and simulation. Therein, we limit the discussion mainly to DL concepts. Thereafter, approaches and concepts specific to individual LTP problems in plasma physics, plasma chemistry, and plasma-surface interactions are addressed.
@@ 3.1 ML methodology
An important variant of RNNs are LSTM ANNs suggested by Schmidhuber and co-workers.[152, 153] While similar in terms of recurrent feedback, their success is largely enabled by memory cells and gate units instead of basic neurons. Each memory cell with a self-connection includes gate units to control the information flow into and out of the cell, as well as of the recurrent feedback. As such, this may be attributed to the concepts of learning and forgetting,[153] and allows for a substantially improved learning efficiency and robustness. LSTM networks have been applied in fluid dynamics for the prediction of the temporal dynamics of turbulent flow through channels based on direct numerical simulation (DNS) data.[154] They have further been used for the prediction of disruption in magnetic confinement fusion plasmas.[155, 156]
@@
Moreover, in a study on a hierarchical approach to multiscale data-driven modeling, LSTM networks have been included for comparison.[157] Another fundamentally related variant of RNNs that was included in this study are echo-state networks (ESN).[151, 158] Therein, an RNN network reservoir is initially set up. A number of inputs are fed into the RNN reservoir, while a number of outputs are extracted from the reservoir. In contrast to ordinary RNN networks, which comprise of learnable weight parameters at every neural node, only the subset of output weight parameters is adjusted during training. Consequently, a manifold of complex nonlinear system dynamics may inherently be evolved by the network, whereas only a small number of output weight parameters capture the desired system dynamics to be described by the data-driven model. To this end, a physics-informed ESN architecture has been investigated for the description of chaotic systems.[159] For the chaotic dynamics of a Lorenz system and a Charney–DeVore system, robust training was achieved with an extended prediction time horizon of approximately two Lyapunov times.
@@
Following the physics-informed learning paradigm,[33] the amount and the variants of approaches exceeds the frame of what can be covered here. The interested reader is referred to an extensive review on PINNs in scientific ML by Cuomo et al. (which also partially covers RNNs and its variants).[160] Another approach to physics-informed ML extending this idea is based on deep neural operator networks (DeepONets).[161] With the reasoning that ANNs can be considered universal operator approximators (similar to the property that ANNs are universal function approximators), these networks comprise of a trunk network and a (single or stack of) branch network(s). The branch part takes as input discrete sensor samples of a function . The trunk part takes as input the quantity . The output of both sub-networks is combined to provide a functional representation . Following an offline training for a variety of examples, Lu et al. have demonstrated online inference performance significantly improving on other reference ANNs (among others LSTM networks). Different variants and extension taking advantage of Fourier and Laplace transformations as nonlinear kernels have further been proposed.[162, 163, 164]
@@
Reduced-order models have previously been used in LTP modeling and simulation as exemplified above. The particular notion of sparse identification of nonlinear dynamics (SINDy),[166, 167] has been demonstrated to provide interpretable and generalizable nonlinear differential equations of reduced order which effectively capture the dynamics of the full model. This is achieved by identifying the relevant system dynamics from a library of potential functional forms by means of a sparse regression. This has been applied to space and astrophysical plasmas based on PIC simulations by Alves and Fiuza demonstrating the potential of recovering the fundamental dynamics of the high-fidelity kinetic plasma simulations.[168] It may be straightforwardly applicable to LTP modeling and simulation data to capture their nonlinear dynamics, albeit a complex chemistry may necessitate further adaptation. Sparse regression has similarly been applied by Kaptanoglu and co-workers in the context of magnetic confinement fusion plasmas for the optimization of magnetic fields in stellarator fusion plasmas.[169, 170] Moreover, MPC based on such ROMs has been suggested with limited availability of training data.[171] An aspect that has been applied using time-dependent simulations of the tokamak plasma boundary.[172, 170] It may be specifically relevant also in the context of LTP processing with its diverse requirements for precise process control.
@@
In contrast to classical supervised learning schemes, alternative approaches may be relevant also due to limited availability of training data. In a Bayesian setting, provided with an estimate of the prediction uncertainty, active learning has been exploited for efficient parameter space exploration.[173] While such uncertainty metric is straightforward in the context of GPR modeling,[22] similar techniques have been adapted to MD simulations and active learning of corresponding ML surrogate models.[174]
Transfer learning enables an alternative route to ML in the small-data limit.[175] The fundamental idea relates to the property of generalization that data-driven models aim pursue. Given an initial problem task with abundant data, a ML model may be robustly trained. The pretrained model may be subsequently retrained on a related, yet different small-data task, whereas just a minor subset of weight parameters are allowed to be adjusted. Hence, the robustness and generalized prediction of the initial model may be successfully transferred to the more challenging learning task. This approach has been successfully applied for instance in the design optimization in inertial confinement fusion.[176]
@@
As surveyed by Gilpin et al.[177], a multitude of explainable AI (XAI) concepts have been proposed for the interpretation and explanation of data-driven models. As an example, local interpretable model-agnostic explanations (LIME) has been suggested for identifying interpretable models that locally approximate the nonlinear behavior of any underlying model.[178] While this is possibly an underexplored aspect in the context of LTP, an example where LIME relates to the data-driven classification and interpretation of OES experiments of plasma in aqueous solution.[179] Further works by Lundberg et al.[180] has included LIME and other methods in an extension Shapley additive explanations (SHAP). It is similarly based on local explanation models, but defines additive feature attribution methods to unify several of the included interpretable methods. There is arguably no mathematical proof that physics-based and physics-informed ML, and corresponding models capture exactly the desired dynamics (and only those). LTP modeling and simulation has got its inherent foundation in a rigorous mathematical description and numerical solution. XAI methods may not only foster advances in data-driven approaches to the field, but strongly contribute to the general acceptance in LTP science and technology.

### v2

@@ 2.3 Plasma-surface interaction
Krüger et al.[125] have conceptualized a PSI ML surrogate model for the Ar ion bombardment of a TiAl composite target as indicated in Figure 3. The ML model, specifically an MLP, was meant to map the incident ions’ energy distributions to the energy angular distributions (EADs) of sputtered particles. It was trained with noisy data from TRIDYN simulations (poor statistics), but demonstrated to be capable of predicting smooth EADs, which were almost congruent to reference TRIDYN simulations with high statistical quality. Gergs et al.[126] have continued those efforts by considering TiAl composite targets with varying compositions, introducing the stoichiometry as a basic system state to the model. A similar data aggregation scheme based on TRIDYN simulations was executed. The ML model was replaced by a combination of a convolutional -variational autoencoder (-VAE)[127, 128, 129] and CNN, defining a reduced dimensional (latent) space for the sputtered particles’ EADs and mapping of the incident ion energy distributions as well as TiAl target composition to this particular latent space, respectively. Predictions and reference TRIDYN simulations agreed with each other, even though the assembled PSI ML surrogate model had got merely 0.39 % of the MLP’s[125] degrees of freedom and, hence, a significantly reduced likelihood for overfitting. Kim et al.[130] have pursued a similar but more advanced approach for the sputtering of fluorinated Si surfaces due to Ar ion bombardment. MD instead of TRIDYN simulations were conducted to obtain information on the emitted (i.e., sputtered, reflected) particles’ EADs as a function of the surface F coverage as well as Ar ions’ incident angle (but not kinetic energy). The aggregated EADs were used for a deep learning-based ROM (DL-ROM), comparing the performance of different ML architectures (i.e., autoencoder[27] (AE), denoising AE[131], VAE, conditional VAE (CVAE)[132]) with each other. The CVAE was found to best its competitors. The PSI ML surrogate model was completed by introducing an additional MLP, which mapped the input information (i.e., surface F coverage, Ar ions’ incident angle) to the latent space established by the DL-ROM, enabling resource efficient predictions with atomic, MD fidelity. Gergs et al.[133] have performed hybrid MD/time-stamped force-bias Monte Carlo (tfMC)[134, 135, 136] simulations to setup a PSI ML surrogate model but treated sputtering of Al targets due to Ar ion bombardment and growth of Al thin films at the substrate as generalized wall interactions. The incident particle flux composition as well as Ar ion energy was varied to setup the data set.[137] Two CVAEs were stacked to setup a physics-separating artificial neural network (PSNN), separating the ion bombardment induced damage from the damage to surface state (e.g., composition, mass density) translation. This network architecture has been argued to allow for training with experimental and simulation data in spite of an inherent discrepancy regarding the information accessibility (damage cannot be measured by similar means). The predictions mitigated the noise but maintained key physical features (e.g., threshold for Ar clustering, reflection probability) for small particle doses. In a follow-up work, [138] the authors have shared a methodology that enabled predictions for particle emission as well as growth during the reactive sputter deposition of AlN thin films in Ar/N plasmas on experimental timescales, that is minutes to hours. First, the MD interaction potential was revised with an evolution strategy implemented in the genetic algorithm-based reactive force field optimizer method (GARFfield)[139] to guarantee accurate PSI simulations.[140] Second, a high-throughput, randomized data generation scheme was pursued to efficiently populate the relevant parameter space with hybrid MD/tfMC simulations.[138] Third, two CVAE were trained to predict either the PSIs or the diffusion processes. They were combined to form a PSNN[133] as PSI ML surrogate model, which can readily be used to complement either plasma simulations or experimental diagnostics. The latter scenario was considered to successfully validate the model. Process times of 45 minutes were predicted within merely 34 GPU hours. In contrast, hybrid MD/tfMC simulations were argued to take more than approximately 8 million CPU years to finish on a comparable case study.
@@ 2.4 Plasma process control and optimization
In a series of contributions, Woelfel et al. have developed MPC schemes for radio-frequency reactive magnetron sputtering fundamentally based on the Berg model.[141] Following the identification of a ROM taking into account the nonlinear process dynamics due to surface poisoning,[142] they have devised a controller for process operation at the transition between poisoned and clean surface modes using the previous reduced model, combined with measurements from multipole resonance probe experiments.[143, 144] An improved reduced model and novel control scheme (also based on the Berg model) has been introduced later, taking into account the plasma state and its dynamical behavior.[145]
@@
The optimization of a plasma etch process has been investigated by Xiao et al. using a MPC scheme based on a physical multi-scale model pairing a 2D fluid description of the discharge and a kinetic Monte Carlo simulation of the surface kinetics. Initially, based on the simulated fluxes at the substrate surface, an RNN surrogate model was set up. Another one was implemented for the surface description, taking into account the stochasticity of the MC model.[59] Subsequently, a similar physical modeling approach has been proposed for a hierarchical surrogate model strategy.[60] Using similar 2D plasma fluid simulation results as for the previous approach, a ROM was devised via proper orthogonal decomposition and Galerkin’s method.[149] This reduced model was argued to capture the intrinsic system dynamics on a low-dimensional representation. It thereafter was used for training another RNN surrogate model for rapid prediction of the plasma discharge dynamics during argon-chlorine etching of silicon. In both works by Xiao et al.[59, 60], MPC schemes have further been developed to control the etching depth and the feature bottom roughness of the process.
@@
With the goal to estimate the internal state of a LTP system in real time despite limited access to observable quantities, a physically-constrained extended Kalman filter (PC-EKF) has been devised by Greve et al.[150] Firstly, the general principle including physical constraints (e.g., prohibiting negative quantities) was assessed at the example of the Lorenz system. The PC-EKF has been shown to closely resemble its chaotic system dynamics. Thereafter, the dynamics of a Hall effect thruster described by a global ionization model were considered. Based on particle balance equations for electrons and ions, the electron temperature was dynamically estimated based on the measured discharge currents. Successful recovery of the electron temperature was demonstrated by the investigation of four different discharge modes subject to varying types of ionization oscillations. The proposed scheme has been argued to benefit from improved robustness with the ability to estimate low-frequency dynamics in the presence of possibly not well resolved high-frequency dynamics (in the model or experiments). In a follow-up study also by Greve et al.,[151] the PC-EKF approach was generalized to more complex Ar and Ar/O chemical reaction sets and global models of ICP sources. Particle balance equations as well as a power balance equation for electrons were taken into account. In an extensive case study, the robustness of the scheme has been demonstrated. Among other cases, a real time estimate of one or multiple unknown state variables based on corresponding limited measurements (e.g., electron absorbed input power based on argon ion density) has been shown, including pulsed operation.
@@ 3 Advances
The need for LTP and plasma chemistry databases have been previously highlighted (see section 2 and referenced literature).[9, 10, 63] Many data-driven advances in modeling and simulation of LTP science and technology may be enabled only by a strategic data generation and acquisition initiative. This is particularly the case since data-driven techniques intrinsically require significant amounts of training data, which need to efficiently explore the relevant parameter space. While there is enormous potential in such data generation and exploitation, we reason that this is a matter of scientific practice and awareness but not a technical obstacle, to a substantial extent. In the following, in contrast, we focus on technical aspects that are hypothesized to have the potential to advance the field, proceeding from ML methods to LTP disciplines. We initially discuss ML concepts and methodology of general relevance to LTP modeling and simulation. Therein, we limit the discussion mainly to DL concepts. Thereafter, approaches and concepts specific to individual LTP problems in plasma physics, plasma chemistry, and plasma-surface interactions are addressed.
@@ 3.1 ML methodology
An important variant of RNNs are LSTM networks suggested by Schmidhuber and co-workers.[154, 155] While similar in terms of recurrent feedback, their success is largely enabled by memory cells and gate units instead of basic neurons. Each memory cell with a self-connection includes gate units to control the information flow into and out of the cell, as well as of the recurrent feedback. As such, this may be attributed to the concepts of learning and forgetting,[155] and allows for a substantially improved learning efficiency and robustness. LSTM networks have been applied in fluid dynamics for the prediction of the temporal dynamics of turbulent flow through channels based on direct numerical simulation (DNS) data.[156] They have further been used for the prediction of disruption in magnetic confinement fusion plasmas.[157, 158]
@@
Moreover, in a study on a hierarchical approach to multiscale data-driven modeling, LSTM networks have been included for comparison.[159] Another fundamentally related variant of RNNs that was included in this study are echo-state networks (ESN).[153, 160] Therein, an RNN network reservoir is initially set up. A number of inputs are fed into the RNN reservoir, while a number of outputs are extracted from the reservoir. In contrast to ordinary RNN networks, which comprise of learnable weight parameters at every neural node, only a subset of output weight parameters is adjusted during training. Consequently, a manifold of complex nonlinear system dynamics may inherently be evolved by the network, whereas only a small number of output weight parameters capture the desired system dynamics to be described by the data-driven model. To this end, a physics-informed ESN architecture has been investigated for the description of chaotic systems.[161] For the chaotic dynamics of a Lorenz system and a Charney–DeVore system, robust training was achieved with an extended prediction time horizon of approximately two Lyapunov times.
@@
Following the physics-informed learning paradigm,[33] the amount and the variants of approaches exceeds the frame of what can be covered here. The interested reader is referred to an extensive review on PINNs in scientific ML by Cuomo et al. (which also partially covers RNNs and its variants).[162] It may be noted that the technical implementation of PINNs is subject to ongoing research. This is partially due to the limited performance and convergence characteristics of PINNs in general, and of ordinary ANNs such as an MLP in particular.[52, 163] Another approach to physics-informed ML extending this idea is based on deep neural operator networks (DeepONets).[164] With the reasoning that ANNs can be considered universal operator approximators (similar to the property that ANNs are universal function approximators), these networks comprise of a trunk network and a (single or stack of) branch network(s). The branch part takes as input discrete sensor samples of a function . The trunk part takes as input the quantity . The output of both sub-networks is combined to provide a functional representation . Following an offline training for a variety of examples, Lu et al.[164] have demonstrated online inference performance significantly improving on other reference ANNs (among others LSTM networks). Different variants and extension taking advantage of Fourier and Laplace transformations as nonlinear kernels have further been proposed.[165, 166, 167]
@@
Reduced-order models have previously been used in LTP modeling and simulation as exemplified above. The particular notion of sparse identification of nonlinear dynamics (SINDy),[169, 170] has been demonstrated to provide interpretable and generalizable nonlinear differential equations of reduced order which effectively capture the dynamics of the full model. This is achieved by identifying the relevant system dynamics from a library of potential functional forms by means of a sparse regression. It has been applied to space and astrophysical plasmas based on PIC simulations by Alves and Fiuza demonstrating the potential of recovering the fundamental dynamics of the high-fidelity kinetic plasma simulations.[171] It may be straightforwardly applicable to LTP modeling and simulation data to capture their nonlinear dynamics, albeit a complex chemistry may necessitate further adaptation. Sparse regression has similarly been applied by Kaptanoglu and co-workers in the context of magnetic confinement fusion plasmas for the optimization of magnetic fields in stellarator fusion plasmas.[172, 173] Moreover, MPC based on such ROMs has been suggested with limited availability of training data.[174] An aspect that has been applied using time-dependent simulations of the tokamak plasma boundary.[175, 173] It may be specifically relevant also in the context of LTP processing with its diverse requirements for precise process control.
@@
In contrast to classical supervised learning schemes, alternative approaches may be relevant also due to limited availability of training data. In a Bayesian setting, provided with an estimate of the prediction uncertainty, active learning has been exploited for efficient parameter space exploration.[176] While such an uncertainty metric is straightforward in the context of GPR modeling,[22] similar techniques have been adapted to MD simulations and active learning of corresponding ML surrogate models.[177]
Transfer learning enables an alternative route to ML in the small-data limit.[178] The fundamental idea relates to the property of generalization that data-driven models pursue. Given an initial problem task with abundant data, an ML model may be robustly trained. The pretrained model may be subsequently retrained on a related, yet different small-data task, whereas just a minor subset of weight parameters are allowed to be adjusted. Hence, the robustness and generalized prediction of the initial model may be successfully transferred to the more challenging learning task. This approach has been successfully applied for instance in the design optimization in inertial confinement fusion.[179]
@@
As surveyed by Gilpin et al.[180], a multitude of explainable AI (XAI) concepts have been proposed for the interpretation and explanation of data-driven models. As an example, local interpretable model-agnostic explanations (LIME) has been suggested for identifying interpretable models that locally approximate the nonlinear behavior of any underlying model.[181] While this is possibly an underexplored aspect in the context of LTP, an example where LIME relates to the data-driven classification and interpretation of OES experiments of plasma in aqueous solution has been proposed.[182] Further works by Lundberg et al.[183] has included LIME and other methods in an extension Shapley additive explanations (SHAP). It is similarly based on local explanation models, but defines additive feature attribution methods to unify several of the included interpretable methods. There is arguably no mathematical proof that physics-based and physics-informed ML, and corresponding models capture exactly the desired dynamics (and only those). LTP modeling and simulation has got its inherent foundation in a rigorous mathematical description and numerical solution. XAI methods may not only foster advances in data-driven approaches to the field, but strongly contribute to the general acceptance in LTP science and technology.